One leg of a right triangle is 17 meters longer than the other leg. If the length of the hypotenuse is 25 meters, find the length of each leg.

Answer by jorel1380(3719)

x2+(x+17)2=25x25
x2+x2+34x+289=625
2x2+34x-336=0
x2+17x-168=0
(x+24)(x-7)=0
x=-24 or 7
The length of the legs are 24m. and 7m..